# MILU: A Multi-task Indic Language Understanding Benchmark

### v2

@@ 6 Conclusion
This work has a few limitations. First, we restricted our study to the top 11 languages due to the lack of readily available questions in low-resource languages, which we aim to address in future work. Second, limited computational resources prevented a thorough evaluation of larger models, such as Llama-3.1-70B and Llama-3.1-405B. Third, the scarcity of questions necessitated translating a portion of the dataset. Finally, our evaluation primarily relies on the log-likelihood approach, which may yield different results compared to other established methods, such as generation-based evaluation and chain-of-thought (CoT) prompting.
@@ Appendix A Details about different Exams
Engineering 500 500 500 500 500 234 500 336 100 500 500 4670
Environmental Science 100 499 100 499 128 100 190 99 100 100 179 2094
@@
Geography 179 500 100 498 219 99 169 100 100 175 192 2331
Health and Medicine 100 499 100 499 100 100 112 100 100 116 107 1933
History 375 498 100 496 273 100 269 100 100 275 391 2977
@@
Language Studies 100 500 100 499 129 100 149 100 100 267 174 2218
Law and Ethics 148 500 100 500 157 100 232 100 100 127 160 2224
Literature and Linguistics 111 499 100 499 120 100 113 100 100 247 142 2131
@@
Music and Performing Arts 100 100 100 151 100 100 100 100 100 100 100 1151
Physics 500 500 359 500 500 190 500 288 100 500 500 4437
@@
Religion and Spirituality 100 100 100 248 100 100 100 100 100 100 100 1248
Social Welfare and Development 100 112 100 193 100 100 100 100 100 100 100 1205
Sociology 115 500 100 500 100 100 159 100 100 111 170 2055

### v3

@@ 6 Conclusion
This work has a few limitations. First, we restricted our study to the top 11 languages due to the lack of readily available questions in low-resource languages, which we aim to address in future work. Second, limited computational resources prevented a thorough evaluation of larger models, such as Llama-3.1-70B-Instruct and Llama-3.1-405B. Third, the scarcity of questions necessitated translating a portion of the dataset. Finally, our evaluation primarily relies on the log-likelihood approach, which may yield different results compared to other established methods, such as generation-based evaluation and chain-of-thought (CoT) prompting.
Acknowledgements
We would like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards building datasets, models, tools and other resources for Indian languages. We are also immensely grateful to the volunteers from the AI4Bharat team for their motivation and meticulous efforts in conducting manual audits.
@@ Appendix A Details about different Exams
Engineering 500 500 500 500 500 234 500 336 100 499 500 4669
Environmental Science 100 499 100 499 128 100 190 99 100 99 179 2093
@@
Geography 179 500 100 498 219 99 169 100 100 171 192 2327
Health and Medicine 100 499 100 499 100 100 112 100 100 115 107 1932
History 375 498 100 496 273 100 269 100 100 267 391 2969
@@
Language Studies 100 500 100 391 129 100 149 100 100 210 151 2030
Law and Ethics 148 500 100 499 157 100 232 100 100 127 160 2223
Literature and Linguistics 111 499 100 495 120 100 113 100 100 154 124 2016
@@
Music and Performing Arts 100 100 100 151 100 100 100 100 100 97 100 1148
Physics 500 500 359 500 500 190 500 288 100 499 500 4436
@@
Religion and Spirituality 100 100 100 247 100 100 100 100 100 100 100 1247
Social Welfare and Development 100 112 100 193 100 100 100 100 100 99 100 1204
Sociology 115 500 100 500 100 100 159 100 100 110 169 2053